# Antimatter Annihilation Vertex Reconstruction with Deep Learning for ALPHA-g Radial Time Projection Chamber

### v1

@@ 1.2 Monte Carlo Simulation of Antihydrogen Annihilation
The Monte Carlo simulation of the response of the rTPC is implemented using the GEANT4 toolkit AGOSTINELLI2003250 ; 1610988 ; ALLISON2016186 . Annihilation products are generated at the wall of the apparatus based on the branching ratios bendiscioli1994antinucleon of the proton-antiproton interactions, and the simulation of their trajectories assumes a 1 T magnetic field. A detailed CAD model (including Penning trap electrodes, superconducting coils, vacuum chambers, etc.) is provided to GEANT4 to simulate the passage of particles through the materials poole2012acad . The GEANT4 implementation of the Photo Absorption Ionization (PAI) model APOSTOLAKIS2000597 is employed to simulate the interaction of the charged particles in the drift volume of the detector (gas medium). The ionization drift towards the anode wires is taken into account using a pre-computed table, calculated with the Garfield++ package schindlergarfield++ . The signals on the wires and on the pads are generated due to the movement of the charges in the rTPC, following the charge multiplication process (avalanche) in the vicinity of an anode wire. A signal template, different for wires and pads, is computed using the Ramo theorem sauli2015gaseous , that permits the derivation of the signal induced on electrodes by a single charge. The functional form of the signals takes into account the shaping electronics, and their amplitude is determined using real data.
@@ 1.3 Standard Reconstruction of Antihydrogen Annihilation Vertex
The current standard method of the $\overline{\mathrm{H}}$ vertex reconstruction proceeds in two stages. The first stage reconstructs the three-dimensional position of individual ionization, otherwise known as spacepoints, caused by the interaction of charged particles with the rTPC gas. The signal templates deployed for the Monte Carlo simulation are also used to determine the arrival times of individual electron clusters on the anode wires and on the pads. Since several pads are involved in each avalanche, positional information is extracted by combining neighboring pad hits, and a unique value for the $z$ coordinate is obtained for each electron avalanche. Each wire signal is matched to the nearest groups of pads by their azimuthal position $\phi$ and the drift time. Finally, the inverse pre-computed table, used in the simulation for electron drift, is used to perform the time projection, thus converting the drift time into the radial coordinate $r$. Moreover, since the rTPC is in a 1 T field, it is necessary to correct for the charge displacement in $\phi$, due to the Lorentz force acting in the drift plane, which is orthogonal to the magnetic field. The triplet of numbers $(r,\phi,z)$ uniquely determines a spacepoint, and it is saved in Cartesian coordinates. This first stage of analysis is used in both the conventional vertex reconstruction and the deep learning method presented in this work.
@@
The standard method is thus based on realistic assumptions on the kinematics of the particles under the influence of the electromagnetic fields. Once the individual $\pi^{\pm}$ tracks in the rTPC signals are clearly identified, this method provides an intuitive way to reconstruct the annihilation vertices. However, identification and clustering of tracks is generally challenging due to electronic noise, multiple scattering, $e^{+}e^{-}$ pair-production following neutral pions ($\pi^{0}$) decays, and the variable number of ($\pi^{\pm}$) particles as a result of $\overline{\mathrm{H}}$ annihilation. If a sufficient number of $\pi^{\pm}$ tracks cannot be identified (i.e., at least two), this method cannot reconstruct the positions of the annihilation vertices. Noise and multiple scattering may also degrade the quality of the fit, resulting in increased error in the position of the reconstructed vertices.
@@ 2.1 Deep Learning Approach for Vertex Reconstruction
In the current development, the goal is to predict the position of $\overline{\mathrm{H}}$ annihilation vertices from the 3D input data. Since this can be framed as a regression task, we need to modify the PointNet classification network for regression tasks. For our application, we focus on the reconstruction of the $z$-coordinate of the vertex, since it represents the position along the vertical axis of the rTPC, where the effect of gravity is measured. In what follows, we will discuss the relevant features of the PointNet architecture and the modifications made to achieve our goal.
@@ 2.2 Data Preprocessing
The data must be preprocessed before being fed into PEAR. First, a fixed number of spacepoints in each event must be determined since PointNet requires the input shape of all the events to be consistent. In choosing the number of spacepoints in a given event, it is important to note that, while a larger number of spacepoints would preserve more complete information, it would also slow down the training iterations due to the larger data size. Taking this into account, we choose the number of spacepoints per event to be 800, since around 95% of the events in the training data set contain 800 or fewer spacepoints.If the number of spacepoints in a given event is larger than 800, the points with the smallest anode wire amplitude are discarded, and if the number of spacepoints is smaller than 800, the points with the largest amplitude are duplicated. This is preferable to padding with zero, since such points alter the distribution of spacepoints, thus affecting the quality of vertex position reconstruction.
@@ 2.3 Training and Construction of an Ensemble
between the simulated $z_{\mathrm{Simulated}}$ and predicted $z_{\mathrm{Reconstructed}}$ vertex coordinate during training. This loss combines the advantages of the Mean Absolute Error (MAE) and the Mean Squared Error (MSE) loss functions, which makes the training of the model less sensitive to the outliers in data than the MSE while still providing smoothness near zero pytorch_huber_loss . After conducting a manual hyperparameter search, we determined that setting $\delta=1$ yielded the best performance. This setting means that for residuals with an absolute value smaller than one, the loss behaves quadratically, whereas for larger residuals, it switches to a linear response. While the minimization of the loss was performed using the training dataset, we also tracked the loss on the $\mathrm{Val}_{A}$ set, a portion of the data that is unseen to the model, to ensure that the model generalizes sufficiently well. The training was performed using an NVIDIA A100 GPU with 100 GB memory, available on the Narval cluster managed by the Digital Research Alliance of Canada. The duration of training was approximately 70 hours.
@@
We initialize our deep learning model with three different random seeds: 0, 1, and 2, and train them separately. The optimized models initialized with each of the random seeds form an ensemble, i.e., PEAR. The criteria for the optimization is as follows. Whenever the model exhibited values of ARA lower than the standard (Helix Fit) method, we saved the weights of the model as a checkpoint. Of the saved checkpoints, the one with the smallest ARA value in the $\mathrm{Val}_{A}$ dataset was chosen to form an ensemble. The results presented in the following section were obtained with PEAR constructed as discussed above.
@@ 4 Conclusions and Future Work
In this work, we explored an approach to reconstructing the position of $\overline{\mathrm{H}}$ annihilation vertices in the rTPC used in the ALPHA-g experiment, based on the PointNet deep learning architecture. We focused on the reconstruction of $z$ coordinate of the vertices (position along the vertical axis of the rTPC) using simulated annihilation events.

### v2

@@ 1 Introduction
The method described in this manuscript relies on simulated data. The conventional method, also developed based on simulated data, was successfully deployed in anderson2023observation , certifying an acceptable level of agreement between simulated and experimental datasets. Nevertheless, the current deep learning approach will also be tested with experimental data in future work. The goal of the present work is to lay the groundwork for integrating deep learning into the ALPHA-g experiment’s analysis pipeline.
@@ 1.2 Monte Carlo Simulation of Antihydrogen Annihilation
The Monte Carlo simulation of the response of the rTPC is implemented using the GEANT4 toolkit AGOSTINELLI2003250 ; 1610988 ; ALLISON2016186 . Annihilation products are generated at the wall of the apparatus based on the branching ratios bendiscioli1994antinucleon of the proton-antiproton interactions, and the simulation of their trajectories assumes a 1 T magnetic field. A detailed CAD model (including Penning trap electrodes, superconducting coils, vacuum chambers, etc.) is provided to GEANT4 to simulate the passage of particles through the materials poole2012acad . The GEANT4 implementation of the Photo Absorption Ionization (PAI) model APOSTOLAKIS2000597 is employed to simulate the interaction of the charged particles in the drift volume of the detector (gas medium). The ionization drift towards the anode wires is taken into account using a pre-computed table, calculated with the Garfield++ package schindlergarfield++ . The signals on the wires and on the pads are generated due to the movement of the charges in the rTPC, following the charge multiplication process (avalanche) in the vicinity of an anode wire. A signal template, different for wires and pads, is computed using the Ramo theorem sauli2015gaseous , which permits the derivation of the signal induced on electrodes by a single charge. The functional form of the signals takes into account the shaping electronics, and their amplitude is determined using real data.
@@ 1.3 Standard Reconstruction of Antihydrogen Annihilation Vertex
The current standard method of the $\overline{\mathrm{H}}$ vertex reconstruction proceeds in two stages. The first stage reconstructs the three-dimensional position of individual ionization, otherwise known as spacepoints, caused by the interaction of charged particles with the rTPC gas. The signal templates deployed for the Monte Carlo simulation are also used to determine the arrival times of individual electron clusters on the anode wires and on the pads. Since several pads are involved in each avalanche, positional information is extracted by combining neighbouring pad hits, and a unique value for the $z$ coordinate is obtained for each electron avalanche. Each wire signal is matched to the nearest groups of pads by their azimuthal position $\phi$ and the drift time. Finally, the inverse pre-computed table, used in the simulation for electron drift, is used to perform the time projection, thus converting the drift time into the radial coordinate $r$. Moreover, since the rTPC is in a 1 T field, it is necessary to correct for the charge displacement in $\phi$, due to the Lorentz force acting in the drift plane, which is orthogonal to the magnetic field. The triplet of numbers $(r,\phi,z)$ uniquely determines a spacepoint, and it is saved in Cartesian coordinates. This first stage of analysis is used in both the conventional vertex reconstruction and the deep learning method presented in this work.
@@
The standard method is thus based on realistic assumptions on the kinematics of the particles under the influence of the electromagnetic fields. Once the individual $\pi^{\pm}$ tracks in the rTPC signals are clearly identified, this method provides an intuitive way to reconstruct the annihilation vertices. However, identification and clustering of tracks is generally challenging due to electronic noise, multiple scattering, $e^{+}e^{-}$ pair-production following neutral pions ($\pi^{0}$) decays, and the variable number of ($\pi^{\pm}$) particles as a result of $\overline{\mathrm{H}}$ annihilation. It is worth noting that the sign of the charged pion is not relevant for the reconstruction of the annihilation vertex, and no attempt is made to determine whether the antiproton has annihilated with a proton or a neutron. If a sufficient number of $\pi^{\pm}$ tracks cannot be identified (i.e., at least two), this method cannot reconstruct the positions of the annihilation vertices. Noise, coupled to the active environment of the experimental hall, and multiple scattering may also degrade the quality of the fit, resulting in increased error in the position of the reconstructed vertices.
@@ 2.1 Deep Learning Approach for Vertex Reconstruction
In the current development, the goal is to predict the position of $\overline{\mathrm{H}}$ annihilation vertices from the 3D input data. Since this can be framed as a regression task, we need to modify the PointNet classification network for regression tasks. For our application, we focus on the reconstruction of the $z$-coordinate of the vertex, since it represents the position along the vertical axis of the rTPC, where the effect of gravity is measured. In this work, the $x$ and $y$ coordinates of the vertex position are not reconstructed, as they are not directly relevant for measuring the gravitational acceleration of $\overline{\mathrm{H}}$. The geometry of the rTPC and the orientation of the electric (drift) and magnetic fields are such that the reconstruction problem separates into its orthogonal components: the radial plane ($x-y$ plane), and the axial direction $z$. This implies that the $x-y$ reconstruction cannot be used directly to infer the quality for our method to determine the $z$ coordinate of the vertex. In what follows, we will discuss the relevant features of the PointNet architecture and the modifications made to achieve our goal.
@@ 2.2 Data Preprocessing
The data must be preprocessed before being fed into PEAR. First, a fixed number of spacepoints in each event must be determined since PointNet requires the input shape of all the events to be consistent. In choosing the number of spacepoints in a given event, it is important to note that, while a larger number of spacepoints would preserve more complete information, it would also slow down the training iterations due to the larger data size. Taking this into account, we choose the number of spacepoints per event to be 800, since around 95% of the events in the training data set contain 800 or fewer spacepoints. If the number of spacepoints in a given event is larger than 800, the points with the smallest anode wire amplitude are discarded, and if the number of spacepoints is smaller than 800, points are randomly duplicated until exactly 800 points are obtained. Padding with duplication is indeed the approach adopted in the original PointNet implementation Qi2017 . This is because the duplicated spacepoints will not affect the outcome due to the max-pooling layer. It is preferable to “padding with zeros”, since the point $(0,0,0)$ lies outside the active volume of the rTPC.
@@ 2.3 Training and Construction of an Ensemble
between the simulated $z_{\mathrm{Simulated}}$ and predicted $z_{\mathrm{Reconstructed}}$ vertex coordinate during training. This loss combines the advantages of the Mean Absolute Error (MAE) and the Mean Squared Error (MSE) loss functions, which makes the training of the model less sensitive to the outliers in data than the MSE, while still providing smoothness near zero pytorch_huber_loss . After empirically confirming that Huber loss outperforms MAE and MSE, we focused on the optimization of the hyperparameter $\delta$. As a result of conducting a manual hyperparameter search, we determined that setting $\delta=1$ yielded the best performance. This setting means that for residuals with an absolute value smaller than one, the loss behaves quadratically, whereas for larger residuals, it switches to a linear response. While the minimization of the loss was performed using the training dataset, we also tracked the loss on the $\mathrm{Val}_{A}$ set, a portion of the data that is unseen to the model, to ensure that the model generalizes sufficiently well. The training was performed using an NVIDIA A100 GPU with 100 GB memory, available on the Narval cluster managed by the Digital Research Alliance of Canada. The duration of training was approximately 70 hours.
@@
We initialized our deep learning model with three different random seeds: 0, 1, and 2, and trained them separately. The optimized models initialized with each of the random seeds form an ensemble, i.e., PEAR. Training was carried out for at least 400 epochs, until the overall loss of the training set converges to stable values or stops consistently improving over 200 epochs. The number of training epochs did not exceed 1000. Whenever the model exhibited values of ARA lower than the standard (Helix Fit) method, we saved the weights of the model as a checkpoint. Of the saved checkpoints, the one with the smallest ARA value in the $\mathrm{Val}_{A}$ dataset was chosen to form an ensemble. The results presented in the following section were obtained with PEAR constructed as discussed above.
@@ 4 Conclusions and Future Work
In this work, we explored an approach to reconstructing the position of $\overline{\mathrm{H}}$ annihilation vertices in the rTPC used in the ALPHA-g experiment, based on the PointNet deep learning architecture. Simulated annihilation events were used to train and evaluate the model for the reconstruction of the $z$ coordinate of the vertices (position along the vertical axis of the rTPC), as it is the quantity directly required for the measurement of the gravitational acceleration of $\overline{\mathrm{H}}$.